t(x)=-x²+7x+1 and r(x)=5x²-2x+8,find (t-r)(2).


1
Expert's answer
2021-01-06T19:51:04-0500

"t(x)=-x^2+7x+1" and "r(x)=5x^2-2x+8"

"(t-r)(x)=-x^2+7x+1-(5x^2-2x+8)=-x^2+7x+1-5x^2+2x-8=-6x^2+9x-7"

"(t-r)(2)=-6\cdot 2^2+9\cdot 2-7=-24+18-7=-13"


Answer: "(t-r)(2)=-13."